A hotel bill for N$ 840.00 is divided between 3 workers in the ratio of their salaries. Revone N$ 15 000, Llewellyn N$ 12 500 and Papilon N$ 7 500. How much does Papilon pay?
Answer by addingup(3677)

Divide their salaries by 1000, then add and you get:
15+12.5+7.5 = 35
840/35 = 24
Now:
24*15 =
24*12.5 =
24*7.5 =
Multiply, use your calculator. That's how much each paid. Check your work, add up the products of your multiplications and you should get 840.